Circular No. 7056

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD W. Kunkel, Las Campanas Observatory (LCO); S. Lawrence and A.

Crotts, Columbia University; and P. Bouchet, S. Heathcote, and R. Probst, Cerro Tololo Interamerican Observatory (CTIO), report brightening of the ‘hotspot’ (IAUC 6665, 6710, 6761) on the inner ring around SN 1987A: “Images obtained on Oct. 31 UT at the LCO 2.5-m telescope (+ Hα filter; 0′′.6 FWHM seeing) show a strong concentration of light at p.a. 31^{o} and radius 0′′.59 on the inner edge of the ring. A comparison was made to images taken on 1997 July 10 using the HST’s WFPC2 in the F656N band, which includes Hα and [N II] at 654.8 nm. Hα and [N II] images from Oct. 31 were combined in the same ratio as in F656N, and the point-spread function of the WFPC2 image was matched to that from Oct. 31. Subtraction of these shows a compact spot that now totals 3.7 percent of the Hα ring brightness, up from 1 percent in 1997 Apr. 26 data from HST/STIS, while the ring appears to have faded by ∼ 15 percent in Hα. No other Hα spots are seen elsewhere on the ring, nor any spots in other bands: [N II] at 658.3 nm, [O III] at 500.7 nm, or continuum at 612 nm. Near-infrared images obtained on 1998 Oct. 6 (4-m CTIO Blanco telescope; 0′′.5 FWHM seeing) show that the hot spot has also brightened considerably in the He I 1083.0nm line, and the spot now has 27 percent of the total brightness of the ring at this wavelength, compared to only 15 percent in images obtained with the same equipment on 1997 Nov. 23. The brightness of the ring has remained constant, or only slightly increased (< 5 percent), over this period. Similar images obtained on 1998 Feb. 13 are consistent with a roughly linear increase in the hot spot’s brightness with time. Comparison of the Oct. 1998 Hα and He I images suggests that the hot spot is slightly more compact in He I and that its photocenter is displaced ∼ 0′′.1 eastward along the ring.

“In addition, R. Blum and S. Heathcote (CTIO) report that lowresolution (R ∼ 550), near-infrared (840-2400 nm) spectra obtained on 1998 Nov. 10 with the 4-m telescope show strong lines of He I at 1083 nm and [Fe II] at 1257, 1278, and 1644 nm. Much weaker lines due to [S III] 953.0-nm, [S II] 1035-nm, Paγ 1094-nm, and [Fe II] 1533-nm were also detected. The line at 1278 nm may include a contribution from Paβ; however, the weakness of Paγ suggests that the [Fe II] line dominates. The spatial profile of the He I 1083-nm line is asymmetric and incipiently double-peaked, with the brighter peak being to the northeast (the slit was oriented in p.a. 30^{o}), consistent with the hot-spot-plus-ring structure seen in images. The other lines are too faint to reliably extract spatial information.”

1998 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green